OpenStudy (anonymous):

Find the exact value by using a half-angle identity; tan(7pi/8)

OpenStudy (anonymous):

@Jhannybean What do you think?

OpenStudy (anonymous):

@Luigi0210 I sure could use your brain right about now as well...

OpenStudy (luigi0210):

What is the half-angle formula again?

OpenStudy (anonymous):

@Luigi0210 Um..it's kinda difficult to put on here but I'll try. sin: sin x/2 = +-sqrt1-cos(x)/2

OpenStudy (luigi0210):

isn't it tan?

OpenStudy (anonymous):

cos: cos x/2 = +-sqrt1+cos(x)/2 tan: tan x/2 = 1-cos(x)/sin(x)

OpenStudy (luigi0210):

There you go

OpenStudy (luigi0210):

Well I got \[\sqrt{2}-1\]

OpenStudy (anonymous):

Really? Huh x) I've seen so many fractions today I nearly forgot about the possiblilty of a whole number (if that classifies as one)

OpenStudy (luigi0210):

I might be wrong.. but do you understand how I arrived at that answer?

OpenStudy (anonymous):

Not really, could you explain?

OpenStudy (luigi0210):

and btw I was wrong it's \[1-\sqrt{2}\]

OpenStudy (luigi0210):

and sure hold on

OpenStudy (luigi0210):

\[\tan \frac{ \theta }{ 2 }=\frac{ 1-\cos \theta }{ \sin \theta } \] \[\tan \frac{ (\frac{ 7\Pi }{ 4 }) }{ 2 }=\frac{ 1-\cos \frac{ 7\Pi }{ 4 } }{ \sin \frac{ 7\Pi }{ 4 } }\]

OpenStudy (luigi0210):

can you figure it out from here?

OpenStudy (anonymous):

lol no, I'm horrible at this stuff, I just want to finish my last couple of tests and be done with this class

OpenStudy (luigi0210):

Ah alright well I'm here if you need anything buddy